OpenStudy (anonymous):

Hey peoples! I need Help! Medal and fan for answer! Four more than three times a number is 16. Let n represent the number. Solve for n. Select one: a. -4 b. 9 c. 4 d. 13

OpenStudy (anonymous):

3n+4=16 3n=16-4=12 n=12/3=4

OpenStudy (anonymous):

The answer is 4? Thanks so much!

OpenStudy (anonymous):

ok